solve the system of equations: a)graphically b)numerically x+2y=4,x+y=1
Answer by mananth(16946)

equation1
when x= 0, y= 2 ( 0 , 2 ) 2
when y = 0, x= 4 ( 4 , 0 )
Equation 2
when x= 0, y= 1 ( 0 , 1 ) 1
when y = 0, x= 1 ( 1 , 0 )



x + 2 y = 4 .............1
x + y = 1 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -2
1 x + 2 y = 4
-2 x + -2 y = -2
Add the two equations
-1 x = 2
/ -1
x = -2
plug value of x in (1)
1 x + 2 y = 4
-2 + 2 y = 4
2 y = 4 + 2
2 y = 6
y = 3
m.ananth@hotmail.ca